# Beyond Complete Shapes: A Quantitative Evaluation of 3D Shape Matching Algorithms

### v1

@@ 4 The BeCoS Benchmark
Partial-to-Full (P2F) Source GT DPFM (U) ULRSSM PFM
@@
Partial-to-Partial (P2P) Source GT DPFM (S) SM-COMB GC-PPSM
@@
Figure 4: Cross-dataset and cross-category shape matching results visualized via colour transfer (red denotes unmatched parts). All baselines show sub-optimal performance on all three settings.
@@
Figure 5: We show geodesic errors (AUC x100) for the baselines on different settings: ful-to-full (left), partial-to-full (middle), partial-to-partial (right). For full-to-full matching shapes can easily be normalized, see Tab. 4. The partial-to-partial setting is the most difficult and more realistic case, in which errors are computed for all vertices. Unmatched vertices are set to an infinite geodesic error; matchings that predict wrong overlaps lead to curves that do not reach 100%.
@@ 5 Evaluation
We evaluate different methods on our benchmark for partial shape matching. To showcase the difficulty that partiality poses, we additionally present the full-to-full results, where we see the research state more saturated than in the partial case. We identify two challenges in the partial shape matching domain, that our benchmark includes.
@@ 5.1 Evaluation Metrics
For quantitative evaluation we consider well-established quantitative measures: we use the geodesic error to quantify the correspondence accuracy using the Princeton Protocol [33], as well as the Intersection over Union (IoU) and the F1 Score to quantify the overlap. To account for the partial-to-partial matching paradigm, when computing geodesic errors we set an infinite geodesic error for those unmatched vertices that should be matched, and for those vertices that are matched but should not be matched. Thus, geodesic error will not reach 100% whenever the matching overlap is imperfect. For more details about the metrics we refer to Sec. 9 in the supp. mat.
@@ 5.2 Baselines
We choose a mix of learning-based and axiomatic methods for each of the three settings (full-to-full, partial-to-full and partial-to-partial), including SOTA methods in the supervised, unsupervised and axiomatic case. We show an overview of the methods in table 2. For the partial-to-partial framework does not yet exist an unsupervised approach. Therefore, we include two axiomatic methods instead. One reason that learning-based approaches are not extensively used for partial-to-partial shape matching is the lack of sufficient training data, which we explicitly address in this work. The severe difficulty of partial-to-partial shape matching is manifested by a significant performance degradation compared to other matching settings, see Fig. 5.
@@ 5.3 Results
We show the results on all three settings in Fig. 5. We additionally show the quality of the overlapping region of the partial-to-partial case in Table 3. Qualitative results can be found in Fig. 4.
@@
mIoU ($\uparrow$) 43.29 47.04 49.34
F1 Score($\uparrow$) 55.58 60.72 62.49
@@
Discussion.
While the learning-based methods on two full geometries yield mostly promising results (see Fig. 4 F2F), existing methods struggle with the challenging partial scenarios presented in our benchmark, demonstrating a need for further methodological advances: In the partial-to-partial case DPFM returns noisy overlapping region predictions (see Fig. 4 P2P, third column). Both, SM-COMB and GC-PPSM perform poorly, probably because they rely on informative and reliable feature descriptors for computing the matching energy. Since these are not available for BeCoS, we resort to XYZ coordinates for computing respective energies (see supp. mat. Sec. 8.1 for a justification and analysis of this choice). In the partial-to-full case both axiomatic and unsupervised methods return bad correspondences both qualitatively (see Fig. 4 P2F, third and fifth column) and quantitatively (see Fig. 5 middle). We discuss further in Sec. 5.4, why our benchmark is so challenging for these settings. We believe that our work will contribute to the development of sophisticated and reliable machine learning approaches for these partial shape matching settings.
5.4 Challenges of Partial Shape Matching
Multiple challenges are simultaneously proposed in the partial shape matching problem: partiality, varying scale, non-isometry, different poses, and non-alignment of pairs of shapes. Our benchmark includes them all, which makes it significantly more challenging than existing datasets. In order to separate the effects of partiality, scaling and rotation we show an ablation study on a normalized setting where all full shapes are set to surface area 1 and different rotation experiments. In contrast to existing partial datasets (e.g. CP2P [3] and PARTIALSMAL [24]), BeCoS shape pairs may be of different sizes, such as a cat and an elephant, which challenges methods to tackle scale-invariant correspondence [29]. We present the effects of the scaling challenge in Section 5.4.1. Furthermore, unlike most previous datasets (e.g. SHREC’16 [17] and CP2P [3]), our shapes are not rigidly aligned. Shapes are randomly rotated around the z-axis (with respect to their canonical pose as present in the original dataset), and thus, methods cannot (explicitly or implicitly) leverage near-rigid alignment for correspondence prediction. We demonstrate how random rotations adversely affect the performance of existing SOTA methods in Sec. 5.4.2.
5.4.1 Challenge 1: Scaling of Shapes
We observe different scaling between the datasets aggregated in BeCoS. To account for this, we manually define scaling factors for each dataset such that shapes are scaled according to semantics (e.g., the same species from different datasets have approximately the same size). In contrast to previous datasets, BeCoS thus contains diverse non-isometrically deformed shapes with different scaling factors (e.g. a cat is smaller than an elephant). We denote this setting as the default setting. We compare this default setting with a normalised setting, where the surface area of each shape in the benchmark is normalised to $1$. These results are depicted in Table 4. Full-to-full methods can easily handle the full-to-full variant of BeCoS under the default setting by normalising the surface area to $1$, resulting in similarly scaled shapes. This is non-trivial for the partial-to-full and partial-to-partial settings, as each partial shape represents an unknown percentage of its full counterpart, making the normalisation non-trivial. Using the normalised surface area under partial- settings thus increases performance, especially for the supervised learning-based methods DPFM [3] and GeomFmaps [21]. Whereas the axiomatic and unsupervised methods perform poorly on BeCoS, making it hard to draw a conclusion about the methods’ performance in the normalised and scaled settings.
Methods Default Normalised +/-
F2F GeomFmaps (S) 89.30 90.08 -0.78
DPFM (S) 90.47 93.29 -2.82
DPFM (U) 88.75 89.42 -0.64
ULRSSM (U) 87.40 94.07 -6.67
Smooth Shells (A) 73.23 73.93 -0.7
P2F GeomFmaps (S) 83.12 87.41 -4.92
DPFM (S) 85.56 89.39 -3.83
DPFM (U) 62.04 61.41 +0.63
ULRSSM (U) 60.09 60.73 -0.64
PFM (A) 62.73 63.49 -0.76
P2P DPFM (S) 37.46 41.75 -4.29
SM-COMB (A) 30.11 25.11 +5.0
GC-PPSM (A) 34.37 32.89 +1.48

### v2

@@ 2.1 Datasets
[…]
Partial-to-Full Datasets. In this setting, a partial shape is deformably matched to a shape template with complete geometry. SHREC’16 [cosmo2016shrec], based on the TOSCA [bronstein2008numerical] dataset, was the first benchmark for this problem setting. PFAUST [bracha2023partial] and PFARM [attaiki2021dpfm] were later introduced using a similar approach based on the FAUST [bogo2014faust] and FARM [kirgo2021wavelet] datasets, respectively. SHREC’20 [dyke2020shrec] also contains a subset of full-to-partial correspondence pairs.
Partial-to-Partial Datasets. Partial-to-partial shape matching involves matching two partial shapes with unknown overlap, making it considerably more challenging than partial-to-full matching. In [attaiki2021dpfm] the authors proposed the first benchmark for this setting with CP2P [attaiki2021dpfm] based on the SHREC’16 CUTS [cosmo2016shrec] benchmark. Recently, PSMAL [ehm2024partial] was introduced, extending the SMAL [Zuffi:CVPR:2017] dataset to include partial-to-partial correspondence pairs.
2.2 Methods for Shape Correspondence
We proceed by discussing a selection of shape matching methods and refer to recent surveys [sahillioglu_recent_2020, deng_survey_2022] for a more thorough treatment. While point-cloud matching methods also exist [li2022lepard, bai2020d3feat, huang2021predator, yu2023rotation], we focus on mesh-based methods, which are better aligned with the surface representations emphasised in our dataset. Methods addressing shape correspondence can be categorised into axiomatic and learning-based. Among the axiomatic approaches are discrete optimization methods [bernard2020mina] and spectral methods utilising the powerful framework of functional maps [ovsjanikov2012functional, rodola2017partial, melzi2019zoomout]. However, the performance of the latter class typically degrades as assumptions of isometry are violated, such as in partial correspondence settings.
Full-to-Full Shape Matching.
Spectral methods based on the eigenfunctions of the Laplace-Beltrami Operator (LBO) have been successful in full-to-full matching settings [ovsjanikov2012functional, litany2017deep, donati2020deepGeoMaps, bastian2023hybrid, cao2023unsupervised], yielding a compact solution. In contrast to spectral approaches, discrete optimization formulations have also been introduced for full-to-full matching [bernard2020mina, roetzer2022scalable, roetzer2023fast, roetzer2024spidermatch]. These algorithms often come with guaranteed bounds on optimality, sometimes incorporating geometric consistency constraints.
Partial-to-Full Shape Matching.
Rodola et al. [rodola2017partial] demonstrate that for partial-to-full matching, the functional map has a slanted diagonal structure, which has, for example, been used in [litany2016non] to solve multi-part partial-to-partial correspondence; however, under the assumption that a full template is given. Several methods also operate in the spectral domain in learning-based settings [attaiki2021dpfm, cao2022unsupervised, cao2023unsupervised]. Other works do not use functional map layer [bracha2023partial, bracha2024wormhole] and explicitly tackle the partial-to-full case.
Partial-to-Partial Shape Matching.
Few works thus far explicitly address the partial-to-partial matching setting, which is particularly challenging as the overlapping region between shapes is unknown. As such, the overlapping region must simultaneously be identified and matched under an unknown deformation. DPFM [attaiki2021dpfm] demonstrates that spectral methods are applicable in this domain using an additional attention mechanism, albeit using ground-truth supervision for feature learning. In [windheuser2011geometrically] shape matching is formulated as an integer linear program (ILP) over the space of orientation-preserving diffeomorphisms. SM-COMB [roetzer2022scalable] demonstrated that the latter constitutes a strong prior to guide partial-to-partial matching. However, holes must be synthetically filled to obtain a closed manifold. Recently, GC-PPSM [ehm2024partial] formulated a geometrically consistent partial-to-partial shape matching non-linear integer program. While both SM-COMB and GC-PPSM are compatible with various surface descriptors, they achieve SOTA results only with high-quality pre-trained input features (e.g. from ULRSSM [cao2023unsupervised]), which can be prohibitive when these are not available [roetzer2022scalable, ehm2024partial].
2.3 Feature Descriptors for Shape Correspondence
Many axiomatic and learning-based methods for shape matching rely on surface feature descriptors as guidance for correspondence computation. Feature descriptors, like SHOT [tombari2010unique], can be handcrafted based on extrinsic quantities like XYZ-coordinates or surface normals. Many surface descriptors are sensitive to surface degradation, such as holes, or are not invariant under rotations, as is the case for XYZ coordinates. The common intrinsic heat kernel (HKS) [sun_concise_2009] and wave kernel signatures (WKS) [aubry_wave_2011] are invariant under isometric deformations; however, they are not effective under partiality as the isometry assumption can be heavily violated. On the other hand, methods like DiffusionNet [sharp2022diffusionnet] aim to learn robust surface descriptors from data while still using some features like XYZ coordinates or WKS descriptors as input. Considering that 3D shapes of many datasets are typically oriented in an upright manner [bogo2014faust, anguelov2005scape, bronstein2008numerical, rodola2014dense, dyke2020shrec], or generally positioned in a canonical pose [eisenberger2021neuromorph] this can lead to biased predictions. We discuss these biases in detail in Sec. 5.4. Moreover, learned features require large training datasets and a high-quality (self-)supervision signals which is challenging to achieve under severe partiality. Since obtaining per-vertex feature descriptors directly on 3D shapes is challenging, [wei2016dense] addresses this by establishing correspondences between 3D shapes through correspondences between their 2D rendered images. Recent methods [dutt2024diff3f, decatur20233d, kobayashi2022decomposing, fischer2024nerf] leverage 2D image foundation model features [oquab2023dinov2, rombach2022high] from multiple rendered views of a 3D shape to aggregate per-vertex features. However, such an approach can suffer from significant bias due to missing semantic information in the rendered views.
Figure 1: Left: We use a collection of multiple complete geometry datasets to generate novel shape pairs. Using FaceForm (see supplementary material), we manually annotate cross-category and cross-dataset correspondences (black lines, see Sec. 3.1) between (manually selected) templates (shapes in red boxes). These are then combined with existing intra-dataset correspondences (blue dashed lines) to construct a shape network, through which dense correspondences can be propagated. Right: Given an input shape with full geometry from our shape collection, we first re-mesh the shape, then apply ray casting to simulate a unidirectional 3D scanning procedure to introduce partiality (see Sec. 3.5), and finally apply a deterministic random rotation.
3 Our Cross-Dataset Shape Matching Network
We propose a novel framework enabling the procedural generation of shape matching instances with ground-truth correspondences based on a given collection of 3D shapes. Our framework supports shapes from multiple categories to generate shape matching instances of non-isometric shape pairs. Furthermore, we manually link correspondences between multiple datasets to vastly increase the diversity and representativeness. Our framework provides researchers with the possibility to propagate dense correspondences or custom annotations (e.g. left/right annotations, which we later consider as demonstrative example) through the entire collection of 3D shapes. Our pipeline supports re-meshing while preserving shape correspondences within a collection and thus fosters discretisation-agnostic method development. To overcome the limitations in realism and sample size of existing partial shape benchmarks, we specifically focus on partial shape generation. We use ray casting to generate more realistic partial shapes than existing benchmarks based on synthetic planar cuts and holes. We summarise the main ideas in Fig. 1. A detailed description of our data generation can be found in the supp. mat. Sec. 9. We apply our framework to construct a new large benchmark combining different shape correspondence datasets (Section 4) and show open challenges in the field through an evaluation of existing methods on this benchmark, as detailed in Sec. 5.
3.1 Cross-Dataset/Category Correspondences
Current datasets commonly lack correspondences between different categories, limiting data availability and diversity [bronstein2008numerical, li20214dcomplete]. We address this shortcoming by manually establishing correspondences across categories and across different datasets. To this end, we treat each category for which a cross-category correspondence is not available due to missing annotation as a separate shape group. Then, we select a template shape for each group and manually choose a set of corresponding landmarks on pairs of templates using FaceForm [faceform] (see supp. mat. Sec. 9). In the case of non-isometry, more manual landmarks are needed to guide the deformation, see Fig. 10 in supp. mat. for an example. We then morph one template into another to establish dense point-to-surface correspondences (represented in barycentric coordinates to account for different discretisations). Using the tightly aligned shapes, dense correspondences are obtained by projecting points onto triangles of the other surface. After the templates are in dense correspondence, we propagate them to the entire shape category using the existing intra-category correspondences. Overall, we build a shape network of pairwise dense correspondences that cover all possible shape templates (cross-category and cross-dataset). The detailed shape network is presented on our project page. A colourmap visualization of the correspondences can be seen in Fig. 1.
3.2 Scaling of Shapes
Our framework integrates multiple mesh-based datasets, each with varying shape scales. To harmonise these datasets, we manually define scaling factors for each dataset, ensuring shapes are scaled according to semantic criteria (e.g., the same species across datasets have consistent sizes). Consequently, our framework offers diverse non-isometrically deformed shapes with varying scaling factors (e.g., a cat is smaller than an elephant; see Fig. Beyond Complete Shapes: A Benchmark for Quantitative Evaluation of 3D Shape Surface Matching Algorithms for examples).
3.3 Alignment of Shapes
Many correspondence datasets, specifically partial ones, such as CP2P [attaiki2021dpfm] and SHREC’16 [cosmo2016shrec], keep shapes in the same canonical pose during training and testing. This bias can prevent methods from generalising to shapes under random rigid transformations, see also [attaiki2023generalizable]. As this assumption is often unrealistic in real-world settings, we randomly rotate the shapes around the z-axis (with respect to their canonical pose as present in the original dataset) and thus generate shapes in random upright positions. We show an ablation study about rotation in Sec. 5.4.
3.4 Custom Annotation Propagation
Shape annotations like movable parts labels used in robotics [qiu2025articulate] or part segmentation labels [kalogerakis2010learn, chen2024DAE-Net] are often indispensable for data-driven training or evaluation. Our framework allows for the propagation of any type of information defined on the vertices from a few annotated examples to all shapes in the dataset. By leveraging the shape network, we can find a path to the closest annotated example from each shape, establishing correspondences to a shape that offers ground-truth information. To demonstrate the effectiveness of our method, we choose to propagate left/right annotations from only $12$ manually annotated shapes to the whole dataset (see Figure 2). While in theory, even fewer annotations would suffice, we selected $12$ shapes to reduce propagation error (see the examples in the supplementary material in Figure 8).
Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b)
Figure 2: Our framework allows for the propagation of custom annotations throughout all shapes. We show the case of left/right annotations. (a): Left/Right annotations are manually annotated on the shape considered as template. (b): Left/Right annotations are propagated from the closest template in the shape network (see Sec. 8.2 in supp. mat. for more examples).
3.5 Partiality Generation
Partiality via Ray Casting. Given a shape $\mathcal{X}$, we introduce partiality by simulating a simple unidirectional 3D scanning procedure based on ray casting, as shown in Fig. 1 right. Doing so leads to a variety of different partial shape boundaries and overlapping behaviour (see Fig. Beyond Complete Shapes: A Benchmark for Quantitative Evaluation of 3D Shape Surface Matching Algorithms for examples). To perform ray casting, we first centre and scale each full shape $\mathcal{X}$ to be within a fixed unit bounding box and then choose random camera poses on the unit sphere. Partiality is obtained by casting $R$ rays from the sampled camera pose onto the full shape $\mathcal{X}$. Rays stop upon hitting a triangle. All hit triangles that form the largest connected component (by area) are kept to be part of the partial shape $\hat{\mathcal{X}}$. Finally, each partial shape $\hat{\mathcal{X}}$ is scaled and translated back to its original size and location. To generate a shape pair for partial-to-partial shape matching we get a partial shape $\hat{\mathcal{Y}}$ from a second full shape $\mathcal{Y}$.
Random camera poses could lead to non-existent or prohibitively small overlap in the partial-to-partial matching setting. Therefore, we sample two camera poses with the constraint that their maximum angular disparity in both azimuth and elevation does not exceed $\alpha$ degrees. Through this, we attain an overlap of a certain size to ensure meaningful matching instances. For the overlap computation, we rigidly align the full shapes $\mathcal{X}$ and $\mathcal{Y}$ with Procrustes analysis using the given correspondences to ensure that both shapes are in a similar position i.e. both shapes are up-right and facing the same direction. Then, we perform the previously described partiality generation on both shapes to generate $\hat{\mathcal{X}}$ and $\hat{\mathcal{Y}}$. Subsequently, we quantify the overlap as the percentage of vertices from $\hat{\mathcal{X}}$ that are mapped to $\hat{\mathcal{Y}}$ and vice-versa. If neither of these quantities falls within a predefined range (see Sec. 4 for details), the process proceeds with new random camera poses until the overlap meets the specified criteria, or until a maximum of $m$ iterations is reached. In the latter case, the partial shapes whose overlap is the closest to the desired range are selected (see Sec. 9.3 in supp. mat. for more details).
Flexible Generation of Shape Matching Instances. Existing partial shape datasets provide a precomputed and fixed number of holes/cuts [cosmo2016shrec, bracha2023partial, kirgo2021wavelet, attaiki2021dpfm, dyke2020shrec, ehm2024partial], creating partial shape pairs from identical meshes in varying poses. Instead, our proposed framework can generate partial shapes from a larger and more diverse pool of shapes, allowing for pairing compatible shapes from different categories, potentially originating from different datasets. This creates a more challenging scenario encompassing simultaneous non-isometric deformations and partiality, representing an important step towards more realistic settings. Furthermore, our framework introduces more variety by using different camera poses for ray casting and re-meshing shapes to different resolutions.
3.6 Semantic Missing Parts
By combining multiple categories and datasets, we encounter cases where dense correspondences cannot be reliably established between certain parts of shapes without introducing inaccuracies (e.g., matching a deer’s horn to a horse’s head). While such ambiguities can be managed when considering semantic correspondences [abdelreheem2023zero], they remain problematic for dense correspondences. To address this, for the first time we create large cross-dataset correspondences with semantic missing parts annotations, as shown in Fig. 3. To this end, we modify our pipeline described in Sec. 3.1 by manually removing the semantic missing parts from the shapes and match the rest of the shapes to get the dense correspondences. We show an example in Fig. 10 in supp. mat. This creates a novel objective for shape matching methods that warrants future research.
Refer to caption Refer to caption Refer to caption Refer to caption
Source Target Source Target
Figure 3: For the first time we define semantic missing parts in full shapes. The horse and puma do not possess a human upper body or horns, making these features unmatchable in the centaur and cow, respectively.
4 The BeCoS Benchmark
In addition to the procedural data generation framework, we propose the BeCoS benchmark, which is a specific instantiation of our general framework introduced in Sec. 3. BeCoS contains realistic full and partial shapes from established datasets with cross-category and cross-dataset correspondences. The purpose of BeCoS is to challenge current state-of-the-art shape matching methods, specifically in the partial case, and facilitate progress in the field through a standardised evaluation. To create our benchmark, we apply the pipeline introduced in Sec. 3 on a combination of several existing datasets. We train and evaluate various existing shape correspondence methods on BeCoS (see Sec. 5.2 for an overview), demonstrating the shortcomings of existing methods, particularly for partial-to-partial matching.
Datasets. BeCoS is composed of seven different datasets with four-legged creatures and humanoid shapes encompassing challenging and diverse shape matching scenarios for both isometric and non-isometric deformations: TOSCA [bronstein2008numerical], FAUST [bogo2014faust], SCAPE [anguelov2005scape], KIDS [rodola2014dense], DT4D [magnet2022smooth], SMAL [Zuffi:CVPR:2017] and SHREC’20 [dyke2020shrec]. See Table 1 for more information about the used datasets. From these datasets, we use a total of 2543 shapes (496 are humanoid shapes and 2047 are four-legged shapes) to generate pairwise matching problems. We partition our dataset into predefined training/validation/test splits, ensuring evaluation is performed on unseen categories. To combat dataset imbalance, we oversample the smaller datasets in our training set. The number of shape matching instances in the train/validation/test splits is 10185/137/142. We limit the test set size to easily compare to axiomatic methods, which are often computationally heavy.
Predefined Benchmark Parameters. We choose the following parameters in the data generation pipeline to create a challenging but reasonable benchmark. For re-meshing, we use quadric decimation [garland1997surface] and we fix the range of vertices to a random number between $9000{-}10{,}000$ to facilitate training of current supervised/unsupervised methods and limit memory consumption. Note that re-meshing the same shape multiple times results in different re-meshed versions of the shape due to the random selection of the number of vertices. Our main setting introduces a random rotation around the z-axis, assuming each shape has a natural ‘upright’ direction. The maximum camera angle for the partial-to-partial setting is set to $\alpha=\frac{\pi}{4}$. Similar to [attaiki2021dpfm] and [ehm2024partial], these parameters are chosen to maintain an overlap of $10\%-90\%$, using $m=10$ maximum iterations (see supp. mat. Sec. 9.3 for distribution of overlap). All default parameters to generate the BeCoS benchmark are explicitly stated and explained in details in our implementation repository.
The parameter configuration of BeCoS is chosen to significantly challenge state-of-the-art methods while still being attainable. In the future, if the benchmark is saturated due to the development of more accurate shape matching methods, the customisable nature of our framework enables a straightforward construction of more challenging BeCoS variants (e.g. random rotation around all axes).
Methods Type F2F P2F P2P Input
features
GeomFMaps [donati2020deepGeoMaps] Sup wks, xyz
DPFM [attaiki2021dpfm] Sup/Unsup wks, xyz
ULRSSM [cao2023unsupervised] Unsup wks, xyz
Smooth Shells [eisenberger2020] Axio shot
PFM [rodola2017partial] Axio shot
SM-COMB [roetzer2022scalable] Axio xyz
GC-PPSM  [ehm2024partial] Axio xyz
Table 2: Overview of baselines with their methodology type (Supervised, Unsupervised or Axiomatic), their possible problem statements (F2F: Full-to-Full, P2F: Partial-to-Full and P2P: Partial-to-Partial) and their corresponding input features. All default values are used for all baselines to produce the results shown in this paper.
Full-to-Full (F2F) Source GT GeomFMaps ULRSSM Smooth Shells
[…]
Partial-to-Full (P2F) Source GT DPFM (Unsup) ULRSSM PFM
@@ 4 The BeCoS Benchmark
Partial-to-Partial (P2P) Source GT DPFM (Sup) SM-COMB GC-PPSM
@@
Figure 4: Cross-dataset and cross-category shape matching results visualised via colour transfer (red denotes missing parts). All baselines show sub-optimal performance on all three settings (some examples are highlighted with blue circles). The green square shows the semantic missing parts between the source and target mesh. As GC-PPSM and SM-COMB are solved on very low resolution when upsampling, we do nearest neighbour search in the features space of the matched patches (see [ehm2024partial] for more information), ensuring geometric consistency in low resolution. Both methods show patch-wise inaccuracies due to this low resolution solution.
@@
Figure 5: We show geodesic errors (AUC x100) for the different methods on different settings: F2F (left), P2F (middle), P2P (right). The partial-to-partial setting is the most difficult and more realistic case, in which errors are computed for all vertices. Unmatched vertices are set to an infinite geodesic error; matchings that predict wrong overlaps lead to curves that do not reach 100%.
@@ 5 Evaluation
We evaluate different shape matching methods, specifically showcasing the difficulty posed by partiality. We additionally present numerous ablation studies and further experiments with custom annotations (left/right annotations), showcasing the challenges and possibilities of BeCoS.
@@ 5.1 Evaluation Metrics
For quantitative evaluation, we consider well-established quantitative measures: we use the Geodesic Error to quantify the correspondence accuracy, as well as the Intersection over Union (IoU) and the F1 Score to quantify the overlap. For the ablation study on custom annotation with left/right prediction, we use a Left/Right Accuracy metric.
Geodesic Error. We evaluate the geodesic error of correspondences using the Princeton Protocol [kim2011blended]. This evaluation entails normalising the geodesic error between the ground truth and the computed correspondences by the shape diameter, defined as the square root of the area of the target shape. For partial shapes, we use the square root of the area of their corresponding full shape. To handle the partial-to-partial matching paradigm, we follow [ehm2024partial] and set an infinite geodesic error for unmatched vertices that should be matched and for vertices that are matched but should not be. In contrast to the partial-to-full (P2F) and full-to-full (F2F) settings, where no overlapping region is determined, the overlapping region prediction affects the geodesic error curve for the partial-to-partial setting.
Intersection over Union. To evaluate the quality of the overlapping areas induced by a matching, we consider the Intersection over Union (IoU). To this end, we define $P\in\{0,1\}^{(|V|\times 1)}$ and $G\in\{0,1\}^{(|V|\times 1)}$ as vectors (predicted and ground truth, respectively), indicating which vertices are part of the matching. We quantitatively evaluate the overlapping region in the partial-to-partial setting using the intersection over union (IoU), calculated as $\text{IoU}=\frac{|P\cap G|}{|P\cup G|}$.
F1 Score. To give further insight into the overlap prediction, we use the F1 score. It integrates precision ($\frac{\sum P\cap G}{\sum P}$) and recall ($\frac{\sum P\cap G}{\sum G}$) through the harmonic mean: F1 score = $\frac{2\cdot\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}$. Its possible value spans $[0,1]$ with $1$ indicating a perfect precision and recall.
Left/Right Accuracy. For a shape $\mathcal{X}$ with $N$ vertices, the left/right accuracy is defined as
$acc_{\chi}=\frac{1}{N}\sum_{i\in N}\mathds{1}(\chi_{i}=\chi_{i}^{gt})$ (1)
where $\chi_{i}$ is the left/right prediction for the $i$-th vertex in $\mathcal{X}$, $\mathds{1}$ is the indicator function, $\chi_{i}^{gt}$ is the ground truth of the left/right annotation of the $i$-th vertex in $\mathcal{X}$. We use a similar approach for the partial-to-partial case as when using the geodesic error. For unmatched vertices that should be matched and vertices that are matched but should not be, we set an accuracy of 00.
@@ 5.2 Baselines
We choose a mix of learning-based and axiomatic methods for each of the three settings (full-to-full, partial-to-full and partial-to-partial), including SOTA methods in the supervised, unsupervised and axiomatic case. We train all learning-based methods on BeCoS. We show an overview of the methods in Table 2. For the partial-to-partial problem setting, an unsupervised approach has not yet been proposed in the literature. Therefore, we include two axiomatic methods instead. One reason that learning-based approaches are not extensively used for partial-to-partial shape matching is the lack of sufficient training data, which we explicitly address in this work. The severe difficulty of partial-to-partial shape matching is manifested by a significant performance degradation compared to other matching settings, see Fig. 5.
@@ 5.3 Results
We show the results on full-to-full, partial-to-full and partial-to-partial settings in Fig. 5. We additionally show the quality of the overlapping region of the partial-to-partial case in Table 3. Qualitative results can be found in Fig. 4.
Discussion. While the learning-based methods yield mostly promising results on the full geometries (see Fig. 4 F2F), existing methods struggle with the challenging partial scenarios (P2F and P2P) presented in our benchmark, demonstrating a need for further methodological advances: In the partial-to-partial case DPFM returns noisy overlapping region predictions (see Fig. 4 P2P, third column). Both SM-COMB and GC-PPSM perform poorly, probably because they rely on informative and reliable feature descriptors for computing the matching energy. Since these are not available for BeCoS, we resort to XYZ coordinates for computing respective energies (see supp. mat. Sec. 8.1 for a justification and analysis of this choice). In the partial-to-full case, both axiomatic and unsupervised methods return bad correspondences, both qualitatively (see Fig. 4 P2F, third and fifth column) and quantitatively (see Fig. 5 middle). We discuss further in Sec. 5.4, why our benchmark is so challenging for these settings. We believe that our work will contribute to the development of sophisticated and reliable machine learning approaches for these partial shape matching settings.
@@
mIoU ($\uparrow$) 46.52 46.91 49.24
F1 Score($\uparrow$) 60.28 60.61 62.42
@@
5.4 Ablation Studies
Multiple challenges can be simultaneously present in the shape matching problem: partiality, varying scale, non-isometry, different orientations/poses, and non-alignment of pairs of shapes. Previous datasets, especially partial shape matching datasets, simplify these challenges (e.g. by aligning the shapes). In contrast, we include all the above mentioned challenges in our benchmark, which makes it significantly more challenging.
In order to separate the effects of partiality, scaling and rotation on existing methods we conduct two ablation studies. First, we present the effects of the scaling challenge. Then, we demonstrate how random rotations adversely affect the performance of existing SOTA methods for partial-to-full shape matching.
Methods Default ($\uparrow$) Normalised ($\uparrow$)
F2F GeomFmaps (Sup) 88.72 88.96 (-0.24)
DPFM (Sup) 90.20 91.47 (-1.27)
DPFM (Unsup) 88.75 89.40 (-0.65)
ULRSSM (Unsup) 87.45 94.06 (-6.61)
Smooth Shells (Axio) 73.23 73.93 (-0.7)
P2F GeomFmaps (Sup) 85.35 89.92 (-4.57)
DPFM (Sup) 85.44 90.02 (-4.58)
DPFM (Unsup) 62.72 61.85 (+0.87)
ULRSSM (Unsup) 60.93 61.00 (-0.07)
PFM (Axio) 60.70 62.83 (-2.03)
P2P DPFM (Sup) 38.11 39.58 (-1.47)
SM-COMB (Axio) 30.03 25.09 (+4.94)
GC-PPSM (Axio) 34.29 32.81 (+1.48)